# Three Puzzles with Covariance and Supertranslation Invariance of Angular Momentum Flux (with Solutions)

### 05 Dec 2023 v1

@@ 1 Introduction
The problem of a proper definition of angular momentum and boost charges for asymptotically flat spacetimes in general relativity has existed since it was discovered in [1, 2, 3, 4] that the metrics of spacetimes allowing for gravitational radiation enjoy an infinite dimensional asymptotic symmetry algebra. This is the BMS (Bondi-Metzner Sachs) algebra, which contains the Poincaré algebra as a proper but not normal subalgebra. The Poisson brackets of angular momentum with supertranslations, which are an infinite-dimensional Abelian subalgebra of BMS, do not vanish; consequently the total angular momentum and the flux of angular momentum through null infinity can be changed by a gravitational wave of infinite wavelength. Since such a wave cannot be detected by any finite-size observer, the ambiguity due to BMS seems to preclude a meaningful definition of angular momentum and the other Lorentz charges in general relativity (GR). Problems inherent in defining angular momentum in GR were noticed by R. Penrose as early as 1964 [5]. In any quantum theory of gravity the supertranslation ambiguity also implies that no unique angular momentum can be assigned to the vacuum.
@@
Renewed interest in the symmetries of asymptotically flat 4D spacetime spurred by direct detection of gravitational waves from black hole mergers produced novel, supertranslation-invariant definitions for angular momentum and angular momentum flux as well as for charges and fluxes of the other Lorentz algebra generators. To our knowledge, the first supertranslation invariant formula for angular momentum was given in ref. [6]. In fact, ref. [6] defines the Bondi charge for angular momentum at any retarded time so it also provides a definition of flux. Refs. [7, 8] give an independent definition of supertranslation-invariant Lorentz Bondi charges which, while agreeing with [6] at retarded time $u\rightarrow-\infty$ differ at finite $u$. Refs. [9, 10] provide formulas for supertranslation-invariant Lorentz charges and their fluxes based on a canonical formalism that applies equally well to the quantum theory.
@@
The angular momentum flux given in [6, 9, 10] begins at $O(G^{3})$ in a perturbative expansion in powers of the Newton constant $G$. This is at variance with explicit computation of mechanical angular momentum flux done by several groups with different methods. In particular both [11] using techniques developed in [12] and [13, 14, 15, 16, 17, 18, 19] agree on the presence of a nonvanishing angular momentum flux at $O(G^{2})$. The origin of this discrepancy was traced back in [20] to the difference between the supertranslation frame used in perturbative calculation of gravitational scattering and a “canonical” frame in which the Bondi angular momentum at $u=-\infty$ agrees with the ADM (Arnowitt, Deser, Misner) definition [22]. Ref. [23] employs the results of [20] to define, to all orders in $G$, a supertranslation-invariant angular momentum flux that agrees to $O(G^{2})$ with perturbative calculations and is defined only in term of asymptotic metric data on $\mathscr{I}$.
@@
In this letter we show that the need to define an “intrinsic” angular momentum flux depending on an independently prescribed origin of coordinates is not a problem but rather the feature that allows us to solve the puzzles due to the ambiguity in the choice of $C_{lm}$, $l=0,1$. We being by defining our notations and deriving the key result that supertranslation-invariant definitions of the flux must also be spacetime translation invariant. Next we derives fully covariant formulas for the charges and fluxes defined in refs. [7, 8]. Key to the covariantization is the use of the transformation law of the boundary graviton under boosts, together with covariant and supertranslation-invariant definitions of the center of mass frame, which are used to determine the first two harmonics of the boundary graviton. Finally, we give a physical interpretation to the supertranslation-invariant charges we previously defined, shows that differences among different prescriptions do not manifest themselves until $O(G^{3})$, and shows how current calculations found in the literature implicitly use the same choice of reference frame as our own.
@@ 2 Notations and transformation properties of covariant quantitites
$\Delta C(\Theta)$ $=$ $C(+\infty,\Theta)-C(-\infty,\Theta),\quad\Delta m(\Theta)=m^{+}(\Theta)-m^{-}(\Theta).$ (7)
@@ 3 Lorentz covariant definitions of angular momentum and boost charges and fluxes
$Jy,\quad y=B(\vec{\beta})\bar{y},\quad\bar{y}=\begin{pmatrix}1\\ 0\\ 0\\ 0\end{pmatrix}.$ (17)
A Lorentz transformation $L$ to a frame moving with velocity $\vec{\beta^{\prime}}$ maps $y$ into $y^{\prime}=LB(\vec{\beta})\bar{y}$ This is not a pure boost from the CMRF but the difference is a pure rotation of the CMRF, $B(-\vec{\beta^{\prime}})LB(\vec{\beta})=R$, hence $LB(\vec{\beta})\bar{y}=B(\vec{\beta^{\prime}})R\bar{y}=B(\vec{\beta^{\prime}})\bar{y}$. So, a translation, supertranslation and covariant definition of the Lorentz charges is
@@ 4 Interpretation of the invariant charges and fluxes
So the covariant, supertranslation-invariant flux reduces to the difference of canonical Bondi charge at +$\infty$+ ∞, computed in the frame where the angular metric at $u=+\infty$ is $h_{AB}+O(1/r^{2})$, minus the canonical Bondi charge at $u=-\infty$, computed in the frame where the angular metric at $u=+\infty$ is also $h_{AB}+O(1/r^{2})$. Refs. [20, 21] argue that $h_{AB}+O(1/r^{2})$ is the frame where the Bondi charge reduces to the ADM charge. This identification suggests a very natural interpretation of $\Delta\mathfrak{J}_{Y}$: it is the canonical charge measured after a gravitational scattering in a “round metric” frame minus the initial canonical charge, also measured in a “round metric” frame. The reference frames are fixed by requiring that the initial metric is round before the scattering occurs, then after the scattering has occurred by requiring that the final metric is also round. This procedure erases any gravitational memory –it is on the contrary based on “forgetting” the initial frame-fixing. The interpretation of the flux and its form is essentially the same as in [7, 24, 25] and [26] once the subtleties due to Lorentz covariance and covariant subtraction of the low-$l$ harmonics in $C^{\pm}$ are properly taken into account using eq. (19) for $C^{-}|_{l\leq 1}$ and either prescription A or B for $C^{+}|_{l\leq 1}$.
@@
5 Difference between two consistent definitions Lorentz charges
@@ Acknowledgements
We thank Gabriele Veneziano, Massimiliano Riva and Geoffrey Compère for useful discussions. M.P. is supported in part by NSF grant PHY-2210349 and by the Leverhulme Trust through a Leverhulme Visiting Professorship at Imperial College, London.

### 08 Apr 2024 v2

@@ 1 Introduction
The problem of a proper definition of angular momentum and boost charges for asymptotically flat spacetimes in general relativity has existed since it was discovered in [1, 2, 3, 4] that the metrics of spacetimes allowing for gravitational radiation enjoy an infinite dimensional asymptotic symmetry algebra. This is the BMS (Bondi-Metzner Sachs) algebra, which contains the Poincaré algebra as a proper but not normal subalgebra. The Poisson brackets of angular momentum with supertranslations, which are an infinite-dimensional Abelian subalgebra of BMS, do not vanish; consequently the total angular momentum and the flux of angular momentum through null infinity can be changed by a gravitational wave of infinite wavelength. Since such a wave cannot be detected by any finite-size observer, the ambiguity due to BMS seems to preclude a meaningful definition of angular momentum and the other Lorentz charges in general relativity (GR). Problems inherent in defining angular momentum in GR were noticed by R. Penrose as early as 1964 [5].
The problem is even sharper in quantum gravity because it implies that two states differing by a supertranslation have the same energy, so in particular the vacuum state is infinitely degenerate. This can be seen as follows: denote schematically with $S^{a}$ the generators of supertranslations and with $J^{\mu\nu}$ ($\mu,\nu=0,1,2,3$) the Lorentz generators, then the commutator $[S^{a},J^{\mu\nu}]$ is nonzero so two zero-energy states differing by a supertranslation have different angular momenta. Explicitly,
$J^{ij}|0\rangle=0\Rightarrow J^{ij}[1+c_{a}S^{a}]|0\rangle=[J^{ij},S^{a}]|0\rangle\neq 0,\qquad i,j=1,2,3,$ (1)
so $|0\rangle$ and $[1+c_{a}S^{a}]|0\rangle$ are different (vacuum) states. In well-defined quantum field theories in flat spacetime and in holographic theories of quantum gravity in Anti de Sitter (AdS) space, instead, the vacuum is unique. The apparent non-uniqueness of vacuum in quantum field theories with spontaneously broken symmetries is resolved because the physical Hilbert space decomposes into superselection sectors [6] while an infinity of states at zero energy is excluded in holhographic theories in AdS because the dual CFT has a unique vacuum. Besides, infinite vacuum degeneracy would violate generalizations of the Bekenstein bound [7] such as [8]. The previous observations suggest that the $J^{\mu\nu}$’s appearing in the BMS algebra may not the correct operators to associate to angular momentum and boosts. Other, unambiguous quantities should be found.
Life is easy in Minkowski spacetime because the Lorentz generators can be defined as $J^{\mu\nu}=\int_{\Sigma}n_{\rho}T^{\rho[\mu}x^{\nu]}$, where $\Sigma$ is a complete Cauchy surface with timelike normal $n_{\rho}$. Unfortunately, this formula becomes ill-defined in general relativity and must be substituted by an integral over a 2-surface at infinity, but this is precisely what introduces supertranslation ambiguities.
The problems we have outlined are not merely formal. In numerical general relativity, ambiguities due to supertranslations have been shown to affect the computation of waveforms [9, 10] that are essential tools for detection and study of black hole mergers. The danger of supertranslation-dependent quantities is that they mix the effect of unobservable background noise due to very long waves to the signal due to a given physical process (gravitational scattering, black hole mergers etc.). This danger manifest both at the level of formal definitions and in dealing with numerical simulations that by construction have finite resolution. Finding unambiguous conserved charges and fluxes in general relativity is therefore a serious problem, all the more urgent now after computational advances in general relativity (see e.g. [11]) and observational discoveries [12].
@@
In fact, the first direct detection of gravitational waves from black hole mergers [12] renewed interest in the quest for a better definition of symmetries and conserved charges in asymptotically flat 4D spacetime. Novel supertranslation-invariant definitions for angular momentum and angular momentum flux as well as for charges and fluxes of the other Lorentz algebra generators have recently appeared in the literature. To our knowledge, the first supertranslation invariant formula for angular momentum was given in ref. [13]. In fact, ref. [13] defines the Bondi charge for angular momentum at any retarded time so it also provides a definition of flux. Refs. [14, 15] give an independent definition of supertranslation-invariant Lorentz Bondi charges which, while agreeing with [13] at retarded time $u\rightarrow-\infty$ differ at finite $u$. Refs. [16, 17] provide formulas for supertranslation-invariant Lorentz charges and their fluxes based on a canonical formalism that applies equally well to the quantum theory.
@@
The angular momentum flux given in [13, 16, 17] begins at $O(G^{3})$ in a perturbative expansion in powers of the Newton constant $G$. This is in disagreement with explicit computation of mechanical angular momentum flux done by several groups with different methods. In particular both [18] using techniques developed in [19] and [20, 21, 22, 23, 24, 25, 26] agree on the presence of a nonvanishing angular momentum flux at $O(G^{2})$. The origin of this discrepancy was traced back in [27] to the difference between the supertranslation frame used in perturbative calculation of gravitational scattering and a “canonical” frame in which the Bondi angular momentum at $u=-\infty$ agrees [28] with the ADM (Arnowitt, Deser, Misner) definition [29]. Ref. [30] employs the results of [27] to define, to all orders in $G$, a supertranslation-invariant angular momentum flux that agrees to $O(G^{2})$ with perturbative calculations and is defined only in term of asymptotic metric data on $\mathscr{I}$.
@@
In this letter we show that the need to define an “intrinsic” angular momentum flux depending on an independently prescribed origin of coordinates is not a problem but rather the feature that allows us to solve the puzzles due to the ambiguity in the choice of $C_{lm}$, $l=0,1$ and to the lack of covariance. These are not mere technicalities but foundational problems that have so far precluded a covariant and unambiguous definition of angular momentum in general relativity and have also introduced redundant quantities, $C_{lm}$, $l=0,1$ without apparent physical meaning. Our paper instead shows that they do have a clear meaning; we will show that they define the origin of the space-time coordinate system. Angular momentum depends on the point from which it is computed, but the quantities that determine that point are nowhere to be found among the asymptotic components of the metric. We will show that the “redundant” s- and -p wave components of the electric shear are precisely those “missing” quantities.
We being by defining our notations and deriving the key result that supertranslation-invariant definitions of the flux must also be spacetime translation invariant. Next we derives fully covariant formulas for the charges and fluxes defined in refs. [14, 15]. Key to the covariantization is the use of the transformation law of the boundary graviton under boosts, together with covariant and supertranslation-invariant definitions of the center of mass frame, which are used to determine the first two harmonics of the boundary graviton. Finally, we give a physical interpretation to the supertranslation-invariant charges we previously defined, shows that differences among different prescriptions do not manifest themselves until $O(G^{3})$, and shows how current calculations found in the literature implicitly use the same choice of reference frame as our own.
@@ 2 Notations and transformation properties of covariant quantitites
$\Delta C(\Theta)$ $=$ $C(+\infty,\Theta)-C(-\infty,\Theta),\quad\Delta m(\Theta)=m(+\infty,\Theta)-m(-\infty,\Theta).$ (8)
@@ 3 Lorentz covariant definitions of angular momentum and boost charges and fluxes
$Jy=0,\quad y=B(\vec{\beta})\bar{y},\quad\bar{y}=\begin{pmatrix}1\\ 0\\ 0\\ 0\end{pmatrix}.$ (18)
This definition is covariant because any Lorentz transformation can be decomposed as $L=BR$, with $B$ a pure boost and $R$ a pure rotation. In the CMRF $R\bar{y}=\bar{y}$ so $L\bar{y}=B\bar{y}$. Definition (18) is also unique because Lorentz transformation $L$ to a frame moving with velocity $\vec{\beta^{\prime}}$ maps $y$ into $y^{\prime}=LB(\vec{\beta})\bar{y}$. This is not a pure boost from the CMRF but the difference is a pure rotation of the CMRF, $B(-\vec{\beta^{\prime}})LB(\vec{\beta})=R$, hence $LB(\vec{\beta})\bar{y}=B(\vec{\beta^{\prime}})R\bar{y}=B(\vec{\beta^{\prime}})\bar{y}$. So, a translation, supertranslation and covariant definition of the Lorentz charges is
@@ 4 Interpretation of the invariant charges and fluxes
So the covariant, supertranslation-invariant flux reduces to the difference of canonical Bondi charge at +$\infty$+ ∞, computed in the frame where the angular metric at $u=+\infty$ is $h_{AB}+O(1/r^{2})$, minus the canonical Bondi charge at $u=-\infty$, computed in the frame where the angular metric at $u=+\infty$ is also $h_{AB}+O(1/r^{2})$. Refs. [27, 28] argue that $h_{AB}+O(1/r^{2})$ is the frame where the Bondi charge reduces to the ADM charge. This identification suggests a very natural interpretation of $\Delta\mathfrak{J}_{Y}$: it is the canonical charge measured after a gravitational scattering in a “round metric” frame minus the initial canonical charge, also measured in a “round metric” frame. The reference frames are fixed by requiring that the initial metric is round before the scattering occurs, then after the scattering has occurred by requiring that the final metric is also round. This procedure “forgets” the initial frame-fixing. The interpretation of the flux and its form is essentially the same as in [14, 31, 32] and [33] once the subtleties due to Lorentz covariance and covariant subtraction of the low-$l$ harmonics in $C^{\pm}$ are properly taken into account using eq. (20) for $C^{-}|_{l\leq 1}$ and either prescription A or B for $C^{+}|_{l\leq 1}$. The Lorentz covariant prescription (20) can be used also to covariantize the flux defined in [16]. It differs from $\Delta\mathfrak{J}_{Y}$ by terms proportional to the gravitational memory $C^{+}-C^{-}$. This computation was performed in the CMRF in [31, 32] but the role of gravitational memory in the definition of angular momentum in general realtivity was noticed long before, e.g. in [34]. We will expand on this question in a future publication [35].
@@
5 Difference between two consistent definitions of Lorentz charges
@@ Acknowledgements
We thank Gabriele Veneziano, Massimiliano Riva and Geoffrey Compère for useful discussions and one of the referees for pointing out to us ref. [34]. M.P. is supported in part by NSF grant PHY-2210349 and by the Leverhulme Trust through a Leverhulme Visiting Professorship at Imperial College, London.